# Understanding In-Context Learning with a Pelican Soup Framework

### v1

@@ Abstract
Many existing theoretical analyses of in-context learning for natural language processing are based on latent variable models that leaves gaps between theory and practice. We aim to close these gaps by proposing a theoretical framework, the Pelican Soup Framework. In this framework, we introduce (1) the notion of a common sense knowledge base, (2) a general formalism for natural language classification tasks, and the notion of (3) meaning association. Under this framework, we can establish a $\mathcal{O}(1/T)$ loss bound for in-context learning, where $T$ is the number of example-label pairs in the demonstration. Compared with previous works, our bound reflects the effect of the choice of verbalizers and the effect of instruction tuning. An additional notion of atom concepts makes our framework possible to explain the generalization to tasks unseen in the language model training data. Finally, we propose a toy setup, Calcutec, and a digit addition task that mimics types of distribution shifts a model needs to overcome to perform in-context learning. We also experiment with GPT2-Large on real-world NLP tasks. Our empirical results demonstrate the efficacy of our framework to explain in-context learning.
In-Context Learning, Pelican Soup Framework
@@ 1 Introduction
Large language models (LM) have demonstrated the ability to perform downstream natural language processing (NLP) classification tasks via in-context learning (ICL) (Chowdhery et al., 2022). That is, they can predict the label of an input by conditioning on the demonstration, i.e., a sequence of input-label pairs (Brown et al., 2020), even when the verbalizers (labels present in the demonstration) are semantically irrelevant to the task, e.g., foo/bar instead of negative/positive (Wei et al., 2023). To understand why LMs trained with nonstructural general text can perform ICL, a line of works base their theoretical analysis on the assumption that the data is from a latent variable model (Xie et al., 2022; Zhang et al., 2023; Wang et al., 2023b).
However, the assumptions on a latent variable model in these works leave a gap between theory and practice. Xie et al. (2022) assumes that the general text for training LMs is from a hidden Markov model, which is an oversimplification of natural language. Zhang et al. (2023) propose a more general latent variable model for the data generation process. Although their model is more realistic than a hidden Markov model, their analyses are based on the assumption that there is a data generation process that generates both the general text for LM training and the demonstration for in-context learning. As demonstrations have a format very different from general text, a question to be answered is whether such a generation process really exists. Additionally, considering the diversity of the semantics natural language can express, it is likely that the training data does not cover all possible values in the latent variable’s space. It requires more characterizations on the latent space to discuss the generalization to semantics unseen in the training data.
Besides, previous theoretical analyses do not fully answer some questions either (Xie et al., 2022; Zhang et al., 2023; Hahn & Goyal, 2023). For example, what is the effect of the choice of verbalizers on ICL performance? How does the distribution of downstream task examples affect ICL? How instruction-tuning helps ICL (Chung et al., 2022)?
To answer these questions, we propose a theoretical framework, the Pelican Soup Framework. We introduce the notion of a commonsense knowledge base (KB) and define the data generation process as the process that generates documents consistent with the KB. We also introduce a formalism for NLP classification tasks. This notion and formalism instantiate a data generation process that satisfies the assumption made by Zhang et al. (2023). With slight adaption on the Zhang et al. (2023), we establish a similar bound for ICL loss under our framework. Additionally, because we see the latent variable as the association between the verbalizers and their meaning, our bound can reflect the effect of the choice of verbalizers. It also reflects how instruction tuning affects ICL performance.
If we further assume that all latent meaning can be represented with logic formula over a finite set of atom concepts as suggested by cognitive science theories Fodor (1975, 2008); Piantadosi (2021), we can characterize the generalization to the unseen meaning during inference time. Having this assumption instantiates a setup compatible with the theory for the emergence of complex skills by (Arora & Goyal, 2023), and we can thus apply their analyses on generalization to our framework. We also discuss our results with the results by Hahn & Goyal (2023), where they bound ICL loss with the description length of the ground truth input-label mapping function.
Inevitably, whether our analyses are meaningful depends on how deep models such as transformers (Vaswani et al., 2017) generalize. Specifically, as most of the previous theoretical analyses do, we show that a perfect LM that estimates the next token distribution perfectly according to the training data distribution can do ICL. However, when doing ICL, LMs are conditioned on the demonstration, which has a low probability mass in the training data. LMs can do ICL only if it can generalize well under this distribution shift from training data to ICL demonstrations.
As the generalization of LMs is difficult to analyze theoretically, we conduct experiments to show empirical evidence for our framework. We construct a synthetic dataset, Calcutec (§4), and a synthetic task, the addition task (§5), to mimic the distribution shifts LMs need to overcome. With Calcutec, we show that LMs can do ICL regardless of the distributional shifts and observe the efficacy of chain-of-thoughts. With the addition task, we observe that larger models generalize better under a certain type of distributional shift, which is aligned with the observation about model scaling in the real world Wei et al. (2022). Finally, in §6, we experiment with natural language and show that using pronouns as verbalizers can lead to comparable performance (with GPT-2 Large), which is aligned with our framework where LMs learn ICL ability by modeling meaning association of words.
@@ 2 The Pelican Soup Framework
2.1 Motivation: Pelican Soup
The Pelican Soup game inspires our framework. It is a game involving a puzzle master who has a story in their mind. The game participants aim to recover the story by asking the puzzle master yes/no questions. An observation is that once the participants recover the story, they can answer any questions about the story. Therefore, the story has a similar role as a latent variable defining the input-output mapping, and the yes/no questions are similar to the demonstrations for in-context learning. We include an example in Appendix A.
Given the above observation, we can study in-context learning by considering why humans can solve Pelican Soup riddles. We conjecture that this is because the person who makes the story and the ones who solve the riddle share the same (or similar) commonsense (McCarthy, 1960) about logical relationships among things in this world (Schank & Abelson, 1988). This inspires us to introduce the notion of a commonsense knowledge base in our framework.
@@ 2.2 Training Data Distribution
Our theory framework is based on the interrelations between the semantics of sentences. As suggested by the long history of linguistics studies that aim at representing sentence semantics with formal languages (Carnap et al., 1968; Bresnan & Bresnan, 1982; Steedman, 1987, 1996; Sag et al., 1999), such as first-order logic (Frege et al., 1879; Peirce, 1883), it is reasonable to make a general assumption on language:
Assumption 2.1 (Sentence).
We assume that a sentence in a language is a sequence of words such that humans can determine whether one sentence entails another sentence or whether two sentences contradict each other.
Motivated by the Pelican Soup game and the early formulation of AI (McCarthy, 1960; Croft, 1993), we introduce the notion of a commonsense knowledge base and use it to restrict the possible combination of sentences that can co-occur in a paragraph:
Assumption 2.2 (Knowledge base).
There is a knowledge base $\mathrm{KB}$, which is a set of sentences, such that any paragraph with non-zero probability mass is a set of sentences such that for any two sentences $x_{1},x_{2}$ in the paragraph, $x_{1}\wedge\mathrm{KB}$ does not contradict with $x_{2}$.
It is natural to assume the existence of this knowledge base. This knowledge base may include statements describing interrelations among the concepts represented by words, as cognitive psychologists have suggested word semantics are based on these interrelations (Siskind, 1996; Murphy, 2004). For example, the knowledge base may contain a rule: “If X moved, then the position of X changed” because the semantic of the verb “move” is embodied in its implication on “position”. Based on this knowledge base, humans can do step-by-step reasoning with language. Humans also avoid writing paragraphs that include contradictory sentences.
We also observe that language allows us to arbitrarily associate concepts to an entity or a pronoun, which we refer to as reference. For example, when the pronoun “she” or the human name “Emily” is present in a paragraph, it is associated with a person that can be described with a set of sentences. The usage of the reference is dependent on the meaning it is associated with and is consistent within its context. For example, if “she” is associated with the sentence “a person who has a house”, then the sentence “she has no property” will have 0 probability mass. Moreover, when we want to refer to “the person who has a house” instead of repeating the sentence again, we use “she” for an abbreviation (when there is no ambiguity).
We can see abstract adjectives such as “good” and “bad” as references too, as their meaning also depends on the context. However, their meaning may not be as variable as pronouns. We reflect this difference with a prior distribution for the meaning a reference is associated with.
For simplicity, we only consider single-word references and assume that such association is consistent throughout a document. We write down an assumption formally:
Assumption 2.3 (Reference).
A set of words $\Gamma\subset\Sigma$ serve as references in the language. For every document in the training data, some $r\in\Gamma$ can be associated with a meaning represented as a set of sentences $Z_{r}$, which follows a prior distribution $\Pr(Z_{r})$. Any $z\in Z_{r}$, $z$ present in the document can be replaced by $r$.
Finally, we assume a document is composed of a set of paragraphs where some references are present:
@@
A document is a concatenation of paragraphs separated with a delimiter $d$ (e.g. a blank line). Also, there is a set of reference $S\subset\Gamma$ such that in each of the paragraphs, a $r\in S$ is present.
2.3 A Formalism for NLP Classification Tasks
@@
With the commonsense knowledge base $\mathrm{KB}$ defined in Assumption 2.2, we propose a simple formalism: For any objective or prescriptive NLP classification task (Rottger et al., 2022) that classifies an input $x$ to one of $|\mathcal{Y}|$ classes, it can be described with some instructions $u$ and some descriptions $v_{i}$ for $i=1,2,\cdots,|\mathcal{Y}|$ such that $x$ is labeled as $y_{i}$ if and only if
$\mathrm{KB}\wedge t\wedge x\models v_{i}.$ (1)
That is, $x$ belongs to class $y_{i}$ if and only if based on the commonsense rules in $\mathrm{KB}$ and the instruction $t$, $x$ entails $v_{i}$. Note that because we can rewrite Eq. 1 as $\mathrm{KB}\wedge x\models u\to z_{i}$, we can combine the task instructions with the description of each class $v_{i}$ for $i=1,2\cdots,|\mathcal{Y}|$. Therefore, we can represent any NLP classification tasks as $\langle v_{1},v_{2},\cdots,v_{|\mathcal{Y}|}\rangle$.
[…]
Assumption 2.5 (Downstream Task).
NLP classification tasks can be defined with the formulation described above. Namely, given any classification task with classes $\mathcal{Y}$, we can define it with the descriptions $\{v_{y}\}_{y\in\mathcal{Y}}$. We further assume that the descriptions of different classes contradict with each other, so every input belongs to only one class.
2.4 Bounding ICL Loss
@@
Our framework is quite general; by adapting and combining the analyses by Zhang et al. (2023) and Hahn & Goyal (2023), we have the following theorem:
Theorem 2.6 (Average ICL Loss).
Let the description of a classification task be $\{z_{y}\}_{y\in\mathcal{Y}}$ and $z^{*}$ be the association between the task description $\{v_{y}\}_{y\in\mathcal{Y}}$ and the verbalizers $\Gamma_{\mathcal{Y}}$ used for ICL. Let $K$ be the constraints used for decoding, $\dot{g}$ be the event where a document belongs to a certain genera. Let $S_{t}=x_{1},r_{2},d,x_{2},r_{2},d\cdots,x_{t},r_{t},d$, where $r_{t}$ is the verbalizer that is associated with the label of $x_{t}$ and $d$ is the delimiter. We have for any positive integer $T$, the average cross-entropy loss of ICL is bounded as:
$\begin{split}&-\frac{1}{T}\sum_{t=0}^{T}\log\Pr(r_{t}|x_{t},S_{t-1},K)\\ &\leq\begin{aligned} &-\frac{1}{T}\log\Pr(z^{*},\dot{g}|K)\\ &-\frac{1}{T}\sum_{i=1}^{T}\log\Pr(r_{t},d|x_{t},z^{*},\dot{g},S_{t-1},K)\\ &-\frac{1}{T}\sum_{i=1}^{T}\log\frac{\Pr(x_{t}|z^{*},\dot{g},S_{t-1},K)}{\Pr(x_{t}|S_{t-1},K)}\end{aligned}\end{split}$ (2)
@@
When the last two terms on the right-hand side are non-negative, Eq. 2 shows the average cross-entropy loss of ICL converges to 00 in $\mathcal{O}(1/T)$. We discuss the three terms in the right-hand side below.
The second term becomes 00 if we set $K$ as the constraint that the next two tokens of $S_{t-1},x_{t}$ must be a verbalizer and the delimiter for all $t$. This is because Assumption 2.2 ensures that $x_{t}$ does not conflict with $r_{t}$ and Assumption 2.5 ensures that $x_{t}$ conflict with verbalizers other than $r_{t}$.
@@
We then look at the last term. This term is 00 when $x_{t}$ is conditionally independent to $z^{*}$ as assumed by Zhang et al. (2023). However, this may be an over-simplification because, in natural language, the transition from $x_{t}$ to its next token depends on the content of $x_{t}$. This assumption may actually be unnecessary for convergence because $x_{t}$ is an example from a downstream task related to $z^{*}$; it is likely that
$\Pr(x_{t}|z^{*},\dot{g},S_{t-1},K)\geq\Pr(x_{t}|S_{t-1},K),$
@@
which implies that this term is non-negative, and we can thus ignore this term.
More rigorously, we can show the following corollary:
@@
Let $\dot{g}$ represent a set of documents whose paragraphs are conditionally independent to each other given $z^{*}$, i.e. $\Pr(x_{1},d,x_{2},d,\cdots,x_{T},d|z^{*},\dot{g})=\prod_{t=1}^{T}\Pr(x_{t},d|z^{*},\dot{g})$. If the downstream task data distribution $\mathcal{D}_{X}$ follow $\Pr(x|z^{*},\dot{g})$, then we can bound the average ICL cross-entropy loss over the downstream task as:
$\begin{split}&\mathop{\mathbb{E}}_{x_{1},x_{2},\cdots,x_{T}\sim\mathcal{D}_{X}^{T}}\left[-\frac{1}{T}\sum_{t=0}^{T}\log\Pr(r_{t}|x_{t},S_{t-1},K)\right]\\ &\leq-\frac{1}{T}\log\Pr(z^{*},\dot{g}|K).\end{split}$ (3)
The right-hand side of Eq. 3 characterizes the convergence rate and reflects the difficulty of doing ICL. If $z^{*},\dot{g},K$ are independent, then we can see this term is proportional to $\Pr(z^{*})$. This implies that when the association between label description and the verbalizer is uncommon in the training data (e.g., associating “positive” to “This movie is bad.”), doing ICL is more difficult.
Eq. 2 also allows us to analyze the scenario where we do not constrain the next token of $S_{t-1},x_{t}$ to be a verbalizer while decoding. We can replace $\dot{g}$ with $\ddot{g}$ that represents the documents satisfying $\dot{g}$ and the constraint $K$ (i.e., verbalizers always follow $x$). [^1] Although $\ddot{g}$ may seem unnatural, this genere of documents corresponds to the PCFG structure assumed in Hahn & Goyal (2023). This ensures the second term of Eq. 2 to be 00. The cost of using $\ddot{g}$ makes the first term in bound bigger because $\Pr(\ddot{g}|z^{*})\leq\Pr(\dot{g}|z^{*})$. This reflects that doing ICL without constraining the next token is harder.
$\Pr(\ddot{g}|z^{*})$ may also be related to instruction tuning. The training examples for instruction tuning are input-output pairs following some format, such as having an instruction (e.g., “Label the example as positive if …”) at the beginning and a prompt after each input (e.g. “[example]. The sentiment of this comment is”). Because for examples that follow the special format, the next token after the prompt is always a verbalizer, these examples belong to the genera $\ddot{g}$. Having these examples in the training set would increase $\Pr(\ddot{g}|z^{*})$ and thus make the ICL loss bound converge faster. This explains why instruction tuning helps ICL.
2.5 Extending the Framework with Formal Language
Assuming a latent model poses a dilemma: Language can encode various meanings, so assuming that the latent space is finite is not reasonable, or the space must be very large. However, if the latent space is infinite or is very large, it is possible that the limited training data does not cover the whole space. Without any assumption on the latent space (e.g., the relation between the states in the space), it is impossible to discuss the generalization to unseen latent states. Thus, we provide an extension to our framework.
We can address this dilemma by extending our framework with an assumption stronger than Assumption 2.1:

### v2

# Pelican Soup Framework: A Theoretical Framework for Language Model Capabilities

@@ Abstract
In this work, we aim to better understand how pretraining allows LLMs to (1) generalize to unseen instructions and (2) perform in-context learning, even when the verbalizers are irrelevant to the task. To this end, we propose a simple theoretical framework, Pelican Soup, based on the logical consistency of the training data, the notion of “reference-meaning association”, and a simple formalism for natural language processing tasks. Our framework demonstrates how linguistic and psychology studies can inform our understanding of language models and is connected to several other existing theoretical results. As an illustration of the usage of our framework, we derive a bound on in-context learning loss with our framework. Finally, we support our framework with empirical experiments and provide possible future research directions.
Pelican Soup Framework:
A Theoretical Framework for Language Model Capabilities
Ting-Rui Chiang University of Southern California tingruic@usc.edu          Dani Yogatama University of Southern California yogatama@usc.edu
@@ 1 Introduction
Large language models (LLMs) have demonstrated the capability to perform downstream natural language processing (NLP) tasks. By following instructions, LLMs can perform tasks with zero-shot examples, demonstrating their reasoning capability. With some input-output examples provided in the prompt, LLMs can also perform tasks without instructions, which is known as in-context learning (ICL) (Chowdhery et al., 2022). In particular, Brown et al. (2020) show that LLMs can perform ICL for classification tasks even when the verbalizers (labels present in the demonstration) are semantically irrelevant to the task, e.g., foo/bar instead of negative/positive Wei et al. (2023). However, it is unclear how pretraining leads to these capabilities.
To explain how LLMs acquire these capabilities, we propose a simple theoretical framework, the Pelican Soup framework in §2. Our framework is based on some very general assumptions, such as the logical consistency of training documents (Assumption 3.2) and the variable meanings with which a phrase can be associated in different contexts (Assumption 3.3). Our framework also includes a simple formalism for NLP tasks, which helps explain why LMs can follow instructions.
In §4, we show how we can use this framework to analyze LLM’s ICL capability, which mitigates the limitations of previous theoretical analyses. For example, in the first theoretical analyses of ICL, Xie et al. (2022) assumes that the general text for training LMs is from a hidden Markov model (HMM), which may be an oversimplification of natural language. In comparison, our framework does not require this strong assumption.
Our framework offers unique insights compared to existing theories. For example, although the generation process by Zhang et al. (2023) is more general than the HMM assumption by Xie et al. (2022), it lacks groundings in real-world linguistic phenomena. Our framework mitigates this limitation, as it helps us better explain the physical meaning of the terms in the bound on the ICL loss and shows how the terms reflect real-world practices, such as the choice of verbalizers and the distribution of test examples. Compared to the work by Hahn and Goyal (2023), our theory allows us to bound ICL loss without relying assumptions on the grammar that generate the data.
Furthermore, in §5, inspired by the cognitive science theories Fodor (1975, 2008); Piantadosi (2021), early development of artificial intelligence  (Siskind, 1996; Murphy, 2004) and formal linguistics Carnap and others (1968); Bresnan and Bresnan (1982); Steedman (1987, 1996); Sag et al. (1999), we provide an extension of our framework to explain why generalization is possible. The extension also connects our framework to other theoretical results. For example, our extension instantiates the complex skills in the theory by Arora and Goyal (2023). In §6, with an additional assumption, we can relate the ICL bound to the description length of the underlying input-output mapping function as done by Hahn and Goyal (2023).
Our work informs future LLM research directions. Scientifically, we shed light on how properties of the training data contribute to the ICL and the instruction-following capabilities of LLMs. The framework also shows how linguistic and psychology studies can inform our understanding of modern NLP. Practically, we highlight the importance of acquiring knowledge about the interrelation between concepts through pretraining. As shown in previous studies, the language modeling objective is inefficient for knowledge acquisition Allen-Zhu and Li (2023); Chiang et al. (2024). We suggest developing a better pretraining technique is crucial for future NLP development. Our proposed experimental setups can also facilitate future studies on LLMs’ acquisition of the capabilities.
@@ 2 The Pelican Soup Framework
We aim our theoretical framework at explaining why LLMs can perform well on prompts for downstream tasks even though the prompts have a different distribution than the training corpus. Therefore, our framework includes assumptions qualifying the training corpus distribution in (§3.1) and a formalism for NLP tasks (§3.2). Later, we will show how this framework allows us to bound the loss of ICL.
3 Motivation
The Pelican Soup game inspires our framework. It is a game involving a puzzle master who has a story in their mind. The game participants ask the puzzle master questions. The goal of the participants is to recover the story based on the answer to the yes/no questions. An observation is that once the participants recover the story, they can answer any questions about the story. This is similar to ICL, where the model is required to “figure out” the underlying task based on the input-output pairs in the given demonstration. Once the model identifies the underlying task, it will be able to predict the label of any input. Therefore, the story has a similar role as a latent variable defining the task, and the yes/no questions are similar to the demonstrations for ICL. We include an example in §A.
Given the above observation, we can study ICL by considering why humans can solve Pelican Soup riddles. We conjecture that this is because the person who makes the story and the ones who solve the riddle share the same (or similar) commonsense (McCarthy, 1960) about logical relationships among the events in this world (Schank and Abelson, 1988). Based on this shared commonsense, the game participants can eliminate hypothesis contradictory to the information that the asked yes/no questions reveal, which allows the participants to uncover the underlying story. This idea of “shared commonsense” inspires us to introduce the notion of “no ambiguity” and “consistency” in our framework.
@@ 3.1 Training Data Distribution
Our theory framework is based on the interrelations between the semantics of sentences. Thus, we first make a general assumption:
Assumption 3.1 (No ambiguity).
For all pairs of sentences $x_{1},x_{2}$ in a language, we assume humans can determine the relationship between $x_{1}$ and $x_{2}$ is contradiction, entailment or neutral.
Assumption 3.1 allows us to qualify sentences that can cooccur in a paragraph with non-zero probability. That is, a paragraph generally does not contain self-contradictory information:
Assumption 3.2 (Consistency).
Any paragraph with non-zero probability mass does not contain two sentences contradicts to each other.
To show how modeling natural language leads to the ICL capability, we further introduce the notion of expression-meaning association as a latent variable. It reflects the fact that language allows us to associate meanings with the surface form expressions quite freely. For example, when “she” or the human name “Emily” is present in a paragraph, it is associated with a certain person of certain characteristics, which exhibit its meaning.
Meanwhile, the usage of the expression depends on the meaning it is associated with and is consistent within its context. For example, if “she” is associated with the sentence “a person who has a house”, then by Assumption 3.2, the sentence “she has no property” will have 0 probability mass. Moreover, when we want to refer to “the person who has a house” instead of repeating the sentence again, we use “she” as an abbreviation.
For simplicity, we only consider single-word expressions and assume that such association is consistent throughout a document.
Assumption 3.3 (Expression-meaning association).
There is a set of words $\Gamma$ such that for every document in the training data, some $r\in\Gamma$ in the document is associated with a meaning represented as a set of sentences $Z_{r}$ with a prior distribution $\Pr(Z_{r})$. Any $z\in Z_{r}$ present in the document can be replaced with $r$ without breaking the logical consistency of the document.
Adjectives such as “good” and “bad” are also expressions that can be associated with variable meanings, and their meanings also depend on the context. However, the association may not be as variable as pronouns’. We reflect this with a prior distribution for the meaning with which an expression is associated later in our theoretical analysis.
Finally, we assume a document is a set of paragraphs where some expressions in $\Gamma$ are present:
@@
A document is a concatenation of paragraphs containing $r\in\Gamma$ separated with a delimiter $d$ (e.g., a blank line).
Based on the above assumptions, we can define a perfect LM.
Definition 3.5 (Perfect LM).
We define an LM to be perfect if it only generates outputs satisfying the above assumptions.
In the following, we will show that a perfect LM can follow instruction and do ICL.
3.2 A Formalism for NLP Tasks
@@
With Assumption 3.1, we propose a simple formalism: For any objective or prescriptive NLP task (Rottger et al., 2022) that maps an input $x$ to a set of acceptable outputs $Y$, that task can be described with some task instructions $u$ such that $u\wedge x\wedge y$ does not cause a contradiction if and only if $y\in Y$.
For example, if the task is a generation task, such as solving a math word problem, the instruction $u$ may specify the format (e.g., “Think step-by-step and output the answer after ####”) and the input $x$ is a math word problem to be solve. LLMs’ response is an acceptable answer if and only if the response does not cause a contradiction to $u$ and $x$, i.e., it follows the instruction and the reasoning process aligns with the problem $x$. When it is a classification task, the instruction $u$ should contain descriptions about each label $y$. For example, we can formulate the sentiment analysis task over movie reviews as $u=\langle v_{+},v_{-}\rangle=\langle\text{``I like the movie''},\text{``I dislike the movie''}\rangle$.
Under this formalism, it is trivial that perfect LLMs can follow instructions and solve tasks. It is because when the prompt for an LLM is the concatenation of the instruction and the task input, $u;x$, Assumption 3.2 (the consistency assumption) ensures that a perfect LLM only generates $y$ such that $y$ is logically consistent to its prompt $u;x$. Based on our NLP task formulation, if $y$ is logically consistent to $u$ and $x$, then $y$ is in $Y$, meaning that $y$ is an acceptable answer.
Two intricate questions remain, how can LLMs perform ICL? and why is it possible for an LM to generalize to unseen instructions? We discuss these two questions in §4 and §5.
4 Bounding ICL Loss
To see how we can analyze ICL with our framework, we first setup a latent variable model with our framework using Assumption 3.3 (the expression-meaning association assumption). In this latent variable model, the latent variables are the association between the expressions in the document and their meanings. Assumption 3.2 (the consistency assumption) defines the support of the distribution of the latent variable conditioning on a context: Associations that break the logical consistency of the context have zero probability. Assumption 3.2 also implies that, give a prefix, continuations that imply invalid associations have zero probability mass.
We analyzing ICL with this latent variable model by focusing on the latent variable for the association between the verbalizers and the meanings. Given a classification task, the underlying latent variable value $z^{*}$ that produces the demonstration is the association where verbalizers are associated with the class descriptions of the task (following the definition of tasks in §3.2). For example, when “foo” is used to represent the positive class of an sentiment classification task, “foo” should be associated with the description of the positive class, i.e., “I like the movie”. If an LLM can infer this underlying association $z^{*}$ based on the examples in the demostration and output a continuation satisfying Assumption 3.2, then it can predict the correct verbalizer for a test example.
@@ 4 Bounding ICL Loss
Formally, adapting and combining the analyses by Zhang et al. (2023) and Hahn and Goyal (2023), we have the following theorem.
Theorem 4.1 (Average ICL Likelihood).
Denote a sequence of input-output pairs as $S_{t}=x_{1},r_{1},d,x_{2},r_{2},\cdots,x_{t},r_{t},d$, where $r_{i}$ is the correct verbalizer with which the label of $x_{i}$ is associated for $i=1,2,\cdots,t$ and $d$ is the delimiter that separates the examples. Let the description of a task that maps inputs to classes $\mathcal{Y}$ be $\{v_{y}\}_{y\in\mathcal{Y}}$ and $z^{*}$ represents that the task descriptions $\{v_{y}\}_{y\in\mathcal{Y}}$ are associated with the corresponding verbalizers $\{r_{y}\}_{y\in\mathcal{Y}}\subset\Gamma$ used for ICL. We have for any $T\in\mathbf{Z}^{+}$, the average log likelihood that the LM predicts the correct verbalizer $r_{t}$ for $t=1,2,\cdots,T$ is bounded as:
$\begin{split}&\frac{1}{T}\sum_{t=0}^{T}\log\Pr(r_{t}|x_{t},S_{t-1})\\ &\geq\begin{aligned} &\frac{1}{T}\log\Pr(z^{*})\\ &+\frac{1}{T}\sum_{t=1}^{T}\log\Pr(r_{t},d|x_{t},z^{*},S_{t-1})\\ &+\frac{1}{T}\sum_{t=1}^{T}\log\frac{\Pr(x_{t}|z^{*},S_{t-1})}{\Pr(x_{t}|S_{t-1})}\end{aligned}\end{split}$ (1)
@@
When the last two terms on the right-hand side are nonnegative, Eq. 1 shows that the average log likelihood converges to 0 in $\mathcal{O}(1/T)$. We discuss the terms on the right-hand side below.
The second term is 0. This is because Assumption 3.2 ensures that the probability $\Pr(r|x_{t})$ is zero if $r$ is not an acceptable output for $x_{t}$ (i.e., when $r\neq r_{t}$ for classification tasks).
@@
We then look at the last term. This term is 0 when $x_{t}$ is conditionally independent of $z^{*}$ as assumed by Zhang et al. (2023). However, this may be an oversimplification because, in natural language, the transition from $x_{t}$ to its next token depends on the content of $x_{t}$. Fortunately, this assumption may actually be unnecessary for convergence because $x_{t}$ is an example from a downstream task related to $z^{*}$; it is likely that
$\Pr(x_{t}|z^{*},S_{t-1})\geq\Pr(x_{t}|S_{t-1}),$
@@
which implies that this term is non-negative, and we can thus ignore this term. More rigorously, we can show the following corollary.
@@
Assuming $z^{*}$ is the association suggesting that the sub-sequences separated by the delimiter $d$ are independent of each other, namely, $\Pr(x_{1},r_{1},d,x_{2},r_{2},d,\cdots,x_{t},r_{t},d|z^{*})=\prod_{t=1}^{T}\Pr(x_{t},r_{t},d|z^{*})$. If the downstream task data distribution $\mathcal{D}_{X}$ follow $\Pr(x|z^{*})$, then we can bound the average ICL likelihood over the distribution as:
$\begin{split}&\mathop{\mathbb{E}}_{x_{1},x_{2},\cdots,x_{T}\sim\mathcal{D}_{X}^{T}}\left[\frac{1}{T}\sum_{t=0}^{T}\log\Pr(r_{t}|x_{t},S_{t-1})\right]\\ &\geq\frac{1}{T}\log\Pr(z^{*}).\end{split}$ (2)
The right-hand side of Eq. 2 characterizes the convergence rate and reflects the difficulty of doing ICL. It reflects that when the association between label description and verbalizer is uncommon in the training data (e.g., associating “positive” with “This movie is bad.”), doing ICL is more difficult.
Note that we can extend the results to generation tasks. For generation tasks, we usually use a separator (or a short text span) between $x_{t}$ and $r_{t}$. We can see the separator as an expression that can be associated with different meanings, so the latent space for $z$ is the meaning the separator can be associated with and $z^{*}$ is the association between the separator and the task instruction (details in §C).
5 Generalization
In §2, we assume a latent model, which, however, poses a dilemma. Language can encode various meanings. If the sequence length is unconstrained, the corresponding semantic space can even be infinite. However, if the latent space is infinite, the limited training data would not cover the entire latent space. Without any assumption on the latent space (e.g., the relation between the states in the space), it is impossible to discuss the generalization to unseen latent states. Thus, we provide an extension of our theoretical framework.